OpenStudy (anonymous):

5 • 4 − 8 ÷ 2 6 10 0 16

OpenStudy (johnweldon1993):

Order of operations again :) 5 • 4 − 8 ÷ 2 5 • 4 = 20 8 ÷ 2 = 4 20 - 4 = 16

OpenStudy (anonymous):

Which phrase matches the algebraic expression below? 3(x + 7) + 10 Three times the difference of x and seven plus ten Three times x minus seven plus ten Three times the sum of x and seven plus ten Three times x plus the sum of seven and ten

OpenStudy (johnweldon1993):

well x + 7 = the sum of x and 7 adding another 10 would be altogether the sum of x and 7 plus 10 And three times that is three times the sum of x and 7 plus 10

OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y if x = −1 and y = −2 0 8 −8 2

OpenStudy (johnweldon1993):

We would just plug them in 4(-1) - 2(-2) What does it come out to ...?

OpenStudy (anonymous):

-8?

OpenStudy (johnweldon1993):

-4 - (-4) remember when you are subtracting a negative...you just change it to positive and adding -4 + 4 = 0

OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y − z, if x = 2, y = 3, and z = −6 −2 8 6 −6

OpenStudy (johnweldon1993):

Okay same thing here 4x - 2y - z plugging in those numbers given 4(2) - 2(3) -(-6) What does THAT come out to?

OpenStudy (anonymous):

6

OpenStudy (johnweldon1993):

4(2) = 8 2(3) = 6 so we have 8 - 6 - (-6) remember subtracting a negative... 8 - 6 + 6 Looks like 8 :)

OpenStudy (anonymous):

\[Select the verbal phrase that corresponds with the given expression. (x – 3)^{2} \] x minus 3 squared x less than 3 squared The quantity 3 minus x squared The difference of x and 3 squared

OpenStudy (johnweldon1993):

Hmm I want to say the first one...but not too sure hang on

OpenStudy (anonymous):

ok

OpenStudy (johnweldon1993):

Yeah the first one is how we all would say it..

OpenStudy (anonymous):

Sarah decides to take her six year old son to the circus. The price for the child's ticket is 4.75 dollars less than the price for the adult's ticket. If you represent the price for the child's ticket using the variable "x", how would you write the algebraic expression for the adult's ticket price?

OpenStudy (johnweldon1993):

Childs ticket = 4.75 less than adults so an adult ticket would be that price....plus 4.75 A = x + 4.75

OpenStudy (anonymous):

Simplify the following expression. 2x − 10y + 3x2 + 7y − 10x 3x2 − 8x − 3y 3x2 − 8x − 17y 3x2 + 8x − 3y 3x2 − 12x + 3y

OpenStudy (johnweldon1993):

I thought you said a couple :P Im assuming that is 3x² so combine like terms 3x² is the only number with a variable squared...so it stays alone 2x - 10x = ...? -10y + 7y = ...?

OpenStudy (anonymous):

3x2 − 8x − 17y

OpenStudy (johnweldon1993):

Not quite...remember it is -10y + 7y You're getting closer to a positive number so this should be 3x² - 8x -3y

OpenStudy (anonymous):

\[ Use the distributive property \to simplify the expression below: \frac{ 1 }{ 2 }(g − 12)\]

OpenStudy (johnweldon1993):

Same thing as before here 1/2 times g = 1/2g and 1/2 times -12 = -6 so we have 1/2g - 6

OpenStudy (anonymous):

\[Evaluate: 32 + (8 − 2) • 4 − \frac{ 6 }{ 3 }\]

OpenStudy (johnweldon1993):

Parenthesis first 8 - 2 = 6 32 + 6 * 4 - 6/3 Now do your multiplication and division 6 * 4 = 24 6/3 = 2 32 + 24 - 2 Left to right 56 - 2 54

OpenStudy (anonymous):

wait i wrote it wrong lol

OpenStudy (anonymous):

\[3^{2}+(8−2)∙4−\frac{ 6 }{ 3 }\]

OpenStudy (johnweldon1993):

Oh okay lol...well again parenthesis 8 - 2 = 6 3² + 6 * 4 - 6/3 then exponents 3² = 9 9 + 6 * 4 - 6/3 now multiplication and division 6 * 4 = 24 6/3 = 2 9 + 24 - 2 left to right 33 - 2 31

OpenStudy (anonymous):

se the distributive property to simplify the expression below: −3(2u − 1) −5u − 3 −6u + 3 −5u + 3 −6u − 3

OpenStudy (johnweldon1993):

-3 times 2u = ...? -3 times -1 = ...? *remember a negative times a negative = a positive*

OpenStudy (anonymous):

−6u + 3

OpenStudy (johnweldon1993):

perfect!

OpenStudy (anonymous):

thank you<33333333333333333333333333

OpenStudy (johnweldon1993):

haha no problem! :)